OpenStudy (anonymous):

Find the vertex of the parabola formed by the following quadratic function: y = -3x2 + 12x - 7 A. (-2,-43) B. (2, 5) C. (-1, -22) D. (1, 2)

OpenStudy (anonymous):

the first coordinate of the vertex of \(y=ax^2+bx+c\) is always \(-\frac{b}{2a}\) you have \(y=-3x^2+12x-7\) so \(a=-3, b=12\) and you get \(-\frac{12}{2\times (-3)}=2\)

OpenStudy (anonymous):

the second coordinate is what you get for \(y\) if \(x=2\)

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

yup